Your a manager in a company that produces rocket ships. Machine A and machine B both produce cockpits and propulsion systems. Machine A ran for 22 hours and produced 3 cockpits and 5 propulsion systems. Machine B ran for 44 hours and produced ed 6 cockpits and 10 propulsion systems. Assume both machines produce cockpits at the same rate and both produce propulsion. Systems at the same rate
Does the system have many solutions
Or yes it does machine A and machine B 4 hours to produce a cockpit and 2 hours to produce. A propulsion system
Or yes it takes machine A and machine B
2 hours to produce a cockpit and 4 hours to produce a propulsion system answer PLEASE

Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

Your a manager in a company that produces rocket ships.
Machine A and machine B both produce cockpits and propulsion systems.
Machine A ran for 22 hours and produced 3 cockpits and 5 propulsion systems.
Machine B ran for 44 hours and produced 6 cockpits and 10 propulsion systems. Assume both machines produce cockpits at the same rate and both produce propulsion Systems at the same rate.
:
let c = hrs to produce one cockpit
let p = hrs to produce one propulsion system
:
Machine A ran for 22 hours and produced 3 cockpits and 5 propulsion systems.
3c + 5p = 22
Machine B ran for 44 hours and produced 6 cockpits and 10 propulsion systems.
6c + 10p = 44
simplify, divide by 2
3c + 5c = 22
These two equations are the same, there are several solutions to c & p, however if c & p have to be integers, the solution is limited to
c = 4, p = 2

Answer by ikleyn(52943)

.

The meaning of this set of words is as dark as a midnight in the forest at no moon.